Iodine is an essential element that maintains proper thyroid function. If you consume a 0.50 mg I2 supplement:

How many I atoms did you consume
Answer #1

Molar mass of I2 = 253.8 g

Number of moles in 0.5 mg = 0.5*10^-3/253.8 = 1.97*10^-6 mol

Number of atoms in 1 mol = 6.022*10^23

Number of atoms in 1.97*10^-6 mol = 1.97*10^-6*6.022*10^23= 1.186*10^18 atoms are consumed